OpenStudy (anonymous):

If x-y=7 and xy=-23/2, find x^2+y^2. Please show all your work.

OpenStudy (anonymous):

the answer is: 26

OpenStudy (anonymous):

Please can you explain

OpenStudy (anonymous):

x-y=7, xy=-23/2 so x^2+y^2=(x-y)^2+2xy=7^2+2(-23/2)=49-23=26

OpenStudy (anonymous):

Thanks